Based on the Pure Expectations Theory of interest rates, if the one-year rate is 4%, and the one-year rate, one year from now, is expected to be 10%, the current two-year rate should be

Select one:

a. 7.0 %

b. 3.0%

c. 6.0%

d. 14.0%
Answer #1

Based on Pure Expectations Theory,

(1 + 2 Yr rate)2 = (1 + 1 Yr Rate) * (1 + 1 Yr Rate 1 Yr from now)

(1 + 2 Yr rate)2 = (1 + 4%) * (1 + 10%)

(1 + 2 Yr rate)2 = 1.04 * 1.10

(1 + 2 Yr rate)2 = 1.144

(1 + 2 Yr rate) = 1.0696

2 Yr Rate = 0.0696

2 Yr Rate = 6.96% = 7.0% (Option a)